OpenStudy (anonymous):

compute the first derivates of the following function y=ln((1+3x)^4/(sq rt(2x+2))

OpenStudy (anonymous):

\[ \frac{ d }{dx } lna= \frac{ a' }{ a }\] So you must use the quotient rule on this one. \[\frac{ (\sqrt{2x+2}*\frac{ d }{ dx }\ln(1+3x))-(\ln(1+3x)*\frac{ d }{ dx }\sqrt{2x+2} }{ 2x+2 }\]

OpenStudy (anonymous):